# Climate impacts of equatorial superrotation

### v1

@@ 4.3 Radiative budget changes and their relation to relative humidity
A second notable finding is that the relative humidity feedback also becomes very small; indeed, the relative humidity changes in the interactive-SST runs (Fig. 11, panel a) are very similar to those for the fixed-SST runs. As a consequence, the feedback becomes strongly dominated by the Planck response. It is also very close to zero in the tropics and subtropics and has a substantial contribution only poleward of $40^{\circ}$.
@@ 4.4 Summary
  • In the deep tropics, between $0^{\circ}$ and $15^{\circ}$, the radiative changes are mainly driven by the forcing (Fig. 10). The circulation changes (specifically the collapse of the MMC, which is not fully compensated by eddy transport changes) result in decreased divergence of the meridional energy transport (Fig. 6), which is balanced by increased OLR. The increased OLR, in turn, is driven by a drying of the atmosphere (Fig. 11) which again finds its source in the circulation changes, as they result in increased divergence of the meridional moisture transport (see Section 6). This behavior is found in both fixed- and interactive-SST runs, although in the latter case a portion of the additional energy convergence can be used to warm the surface and the whole column, increasing its specific humidity at the same time. This feedback occurs essentially at fixed relative humidity.
  • Further from the equator, in the subtropics (between $15^{\circ}$ and $40^{\circ}$), the circulation changes (combination of reduced mean meridional circulation and increased eddy transport) lead to decreased convergence of meridional energy transport (Fig. 6), balanced by decreased OLR. The OLR decrease results from the interplay of humidity changes and thermal structure changes (the atmosphere is moister and slightly cooler, Figs. 8 and 9) which are well summarized by a strong increase in relative humidity (Fig. 11).Again, this relative humidity change is consistent with the circulation changes, and in particular the collapse of the MMC. The interactive-SST runs are quite similar to the fixed-SST runs in that area, i.e. the feedbacks remain relatively small, as the deficit of energy convergence can be compensated radiatively through the humidity increase without affecting surface temperature. Nevertheless, this is less and less the case as we move poleward, and there is a progressive transition to the regime described next.
  • Poleward of $40^{\circ}$, the circulation changes lead to increased convergence of meridional energy transport, driven by a combination of the MMC and the eddy transport (Fig. 6), which provide contributions of comparable magnitudes. Like in the tropics, this is balanced by increased OLR, but the adjustment mechanism is quite different. Here, temperature and specific humidity in the atmosphere change (Figs. 8 and 9) in ways which depend on latitude (warmer and moister in the equatorward part of the region, cooler and drier in the poleward part), with specific humidity not following the Clausius-Clapeyron scaling (i.e, there are substantial relative humidity changes; see Fig. 11). But these changes essentially compensate in the radiative budget, and the balance is reached by uniformly warming the column and the surface, i.e. by the Planck response (because this is not possible in the fixed-SST case, the surplus of energy is absorbed by the surface fluxes).
@@
where $C$ is the heat capacity of the layer where the diffusion process happens, $R$ is the TOA radiative balance (positive downwards), $G$ is the surface radiative balance (positive downwards), $D$ is the diffusivity coefficient and $h$ the generalized enthalpy used in the diffusive closure. Historically, dry static energy was used for the diffusive closure, and the model resulted in a diffusion equation for temperature, with sources and sinks determined by the radiative budget, providing perhaps the simplest climate model with spatial structure (Budyko, 1969; Sellers, 1969; North et al., 1981). More recently it has been argued that such models could also describe finer scale properties of climate, such as the regional response to radiative forcing, for instance in the context of polar amplification, provided the water vapor feedback is taken into account by using MSE as the diffusing quantity (Flannery, 1984; Rose et al., 2014; Merlis and Henry, 2018).
Here we use this moist EBM framework as a way to relate energy budget changes, due to meridional energy transport and radiative budget changes, to surface temperature. In particular, we seek to explain how comparable energy budget changes in the tropics, subtropics, and extratropics result in small, negligible, and large surface temperature changes, respectively. To do so, we solve the 1D diffusion equation for the control and forced runs by prescribing the meridional structure of the radiative budget from the zonally-averaged GCM simulation output, and plot the surface temperature difference on Figure 12 (panel a), for four different levels of complexity:
  • Dry model with constant diffusivity: Temperature profile changes at equilibrium satisfy $\delta R+\partial_{y}D\partial_{y}\delta T=0$ [see (6)]. We use either a uniform diffusivity or a latitude-dependent diffusivity computed empirically from the control run:
    $D=-\left(a^{2}\cos(\phi)\partial_{\phi}\left(T_{\mathrm{s}}+\frac{L_{\mathrm{v}}}{c_{p}}\mathrm{RH}_{\mathrm{s}}q^{\star}(T_{\mathrm{s}})\right)\right)^{-1}\cdot\overline{\left[v\cdot\left(c_{p}T+L_{\mathrm{v}}q+gz\right)\right]},$ (7)
    regularized at the equator (since both the meridional MSE transport and the meridional MSE gradient vanish). In both cases, we use the same diffusivity for the control and superrotating runs. Taking into account, or not, the meridional structure in diffusivity does not affect the results.
@@
  • Moist model with varying relative humidity: Temperature changes satisfy $\delta R+\partial_{y}\left(D\partial_{y}\left(\delta\mathrm{MSE}/c_{p}\right)\right)=0$, where $\delta\mathrm{MSE}=c_{p}\delta T+\delta(\mathrm{RH}q^{\star}(T))$. The relative humidity profiles for the two states are shown in Figure 12 (panel b).
@@
We use the dedalus framework (Burns et al., 2020) to solve for the equilibrium (i.e., steady state) temperature distribution as a boundary value problem, where we constrain the global average of temperature to be the same as in the GCM simulation and assume the meridional flux of energy vanishes at the poles. In the moist case, for the sake of simplicity, we linearize the temperature dependence of the saturation specific humidity, writing $q^{\star}(T)=\alpha T+\beta$.
@@ 6 Moisture transport and precipitation changes
It follows from the water mass budget that the $\mathrm{P-E}$ changes associated with a change in $F_{0}$ can be decomposed into a sum of different contributions, as done for example by Seager et al. (2010); Byrne and O’Gorman (2015) for anthropogenic climate change:
@@
Figure 13: Decomposition of $\mathrm{P-E}$ changes [see Eq. (8)] between the $F_{0}=0$ and $F_{0}=4$ m.s11-1- $-1$ experiments in the interactive- (panel a) and fixed-SST runs (panel b).
As shown in Figure 13, the different components of Eq. (8) behave differently when increasing the forcing. First of all, we note that both in the fixed- and interactive-SST experiments, two components of the decomposition are clearly more important than the others: the mean dynamic component and the eddy component. The two other components (mean thermodynamic and non-linear) are negligible. Note that this is qualitatively different from simulations of anthropogenic climate change where the thermodynamic effect tends to dominate, leading to an approximate Clausius-Clapeyron scaling for $\mathrm{P-E}$ (Held and Soden, 2006), which holds relatively well over ocean regions (Byrne and O’Gorman, 2015).
We start by discussing the interpretation of the two dominant terms in the interactive-SST case (Fig. 13, a). The spatial structure of the total $\mathrm{P-E}$ change mostly matches the structure of the contribution due to MMC changes $\delta\mathrm{MDyn}$. These changes are directly related to the response of the MMC to the imposed tropical torque (Fig. 1) and are consistent with the meridional energy transport changes (Section 44.1): the collapse of the Hadley cells accounts for the negative sign of $\delta\mathrm{MDyn}$ within approximately $10^{\circ}$-$10^{\circ}$ and its positive sign poleward of $15^{\circ}$ in each hemisphere. Similarly, the large shift of the Ferrell cells contributes to the poleward flank of the region of positive moisture convergence changes between $15^{\circ}$-$40^{\circ}$, and explains the negative changes around $50^{\circ}$ as well as the small positive contribution closer to the poles.
@@ 7 Conclusion
To unravel the mechanisms connecting the circulation changes to these temperature and precipitation changes, we have utilized a forcing/feedback analysis using radiative kernels along with decompositions of the energy and moisture transport changes into their various dynamical and thermodynamical components. These analyses reveal that different regions of the planet respond very differently to the dynamical forcing. First, the response in the tropics and subtropics is mostly governed by the moisture changes induced by the collapse of the MMC. This results in large relative humidity changes (decreasing in the tropics and increasing in the subtropics), which allow the atmosphere to compensate radiatively the changes in meridional energy transport related to circulation changes (larger energy convergence in the tropics, smaller in the subtropics). In other words, the radiative budget changes are dominated by the radiative forcing due to relative humidity, and feedbacks are small. Hence, in these regions surface temperature changes are also small. However, precipitation changes are large, and they could affect significantly the regional water cycle and therefore the biosphere. On the other hand, in the mid- to high-latitudes, the effects of moisture transport changes are too weak to compensate the energy transport changes, and the whole atmospheric column as well as the surface must warm to equilibrate the increased horizontal energy flux convergence. In other words, the temperature response is dominated by the Planck feedback. On a more technical note, this work demonstrates how insight can be gained into climate problems where the forcing is dynamic in origin, using various established conceptual tools (radiative forcing/feedback framework, fixed relative humidity ansatz, dry and moist EBMs, dynamic-thermodynamic decomposition of moisture transports).
Future work could investigate the robustness of the results presented here, for instance using different climate models and less idealized setups. In particular, the role of clouds or limited moisture availability at land surfaces are likely to influence the climate responses to forced circulation changes. It also remains to be better understood under which circumstances the climate of the Earth could spontaneously undergo such a transition, if at all, and how the two-way thermodynamic coupling between the processes driving the transition to superrotation and the impacts of superrotation described here would unfold. Finally, if the climate changes described here prove to be robust, a natural question would be whether they can be connected to paleo proxies for warm climates of the past.

### v3

@@ 4.2 Radiative budget changes and their relation to relative humidity
A second notable finding is that the relative humidity feedback is also very small; indeed, the relative humidity changes in the interactive-SST runs (Fig. 8, panel a) are very similar to those for the fixed-SST runs. As a consequence, the feedback becomes strongly dominated by the Planck response. It is also very close to zero in the tropics and subtropics and has a substantial contribution only poleward of $40^{\circ}$.
@@ 4.3 Summary
  • In the deep tropics, between $0^{\circ}$ and $12^{\circ}$ latitude, the radiative changes are mainly driven by the radiative forcing (Fig. 7). The circulation changes (specifically the collapse of the MMC, which is not fully compensated by eddy transport changes) result in decreased divergence of the meridional energy transport (Fig. 6), which is balanced by increased OLR. The increased OLR, in turn, is driven by a drying of the atmosphere (Fig. 8) which again finds its source in the circulation changes, as they result in increased divergence of the meridional moisture transport (see Section 6). This behavior is found in both fixed- and interactive-SST runs, although in the latter case a portion of the additional energy convergence can be used to warm the surface and the whole column, increasing its specific humidity at the same time (see Appendix A). This feedback occurs essentially at fixed relative humidity.
  • Further from the equator, in the subtropics (between $12^{\circ}$ and $40^{\circ}$ latitude), the circulation changes (combination of reduced mean meridional circulation and increased eddy transport) lead to decreased convergence of meridional energy transport (Fig. 6), balanced by decreased OLR. The OLR decrease mostly results from a strong increase in relative humidity (Fig. 8). Again, this relative humidity change is consistent with the circulation changes, and in particular the collapse of the MMC. The interactive-SST runs are quite similar to the fixed-SST runs in that area, i.e. the feedbacks remain relatively small, as the deficit of energy convergence can be compensated radiatively through the humidity increase without affecting surface temperature. Nevertheless, this is less and less the case as we move poleward, and there is a progressive transition to the regime described next.
  • Poleward of $40^{\circ}$, the circulation changes lead to increased convergence of meridional energy transport, driven by a combination of the MMC and the eddy transport (Fig. 6), which provide contributions of comparable magnitudes. Like in the tropics, this is balanced by increased OLR, but the adjustment mechanism is quite different. Although the relative humidity decreases in the equatorward part of the region and increases in the poleward part, the associated radiative response is very small, similar to the lapse rate response. Consequently, the balance is reached by uniformly warming the column and the surface, i.e. by the Planck response (because this is not possible in the fixed-SST case, the surplus of energy is absorbed by the surface fluxes).
@@
where $C$ is the heat capacity of the layer where the diffusion process happens, $R$ is the TOA radiative budget (positive downwards), $G$ is the surface radiative budget (positive downwards), $D$ is the diffusivity coefficient and $h$ the generalized enthalpy used in the diffusive closure. Historically, dry static energy was used for the diffusive closure, and the model resulted in a diffusion equation for temperature, with sources and sinks determined by the radiative budget, providing perhaps the simplest climate model with spatial structure (Budyko1969; Sellers1969; North1981). More recently it has been argued that such models could also describe finer scale properties of climate, such as the regional response to radiative forcing, for instance in the context of polar amplification, provided the water vapor feedback is taken into account by using MSE as the diffusing quantity (Flannery1984; Rose2014; Merlis2018).
Here we use this moist EBM framework as a way to relate energy budget changes, due to meridional energy transport and radiative budget changes, to surface temperature. In particular, we seek to explain how comparable energy budget changes in the tropics, subtropics, and extratropics result in small, negligible, and large surface temperature changes, respectively. To do so, we solve the 1D diffusion equation for the control and forced runs by prescribing the meridional structure of the radiative budget from the zonally-averaged GCM simulation output, and plot the surface temperature difference on Figure 9 (panel a), for four different levels of complexity (below we denote meridional derivatives with $\partial_{y}$):
  • Dry model with constant diffusivity: Temperature profile changes at equilibrium satisfy $\delta R+\partial_{y}\left(D\partial_{y}\delta T\right)=0$ [see (4)]. We use either a uniform diffusivity or a latitude-dependent diffusivity computed empirically from the control run so that the relation
    $\overline{\left[v\cdot\mathrm{MSE}\right]}=D\partial_{y}\overline{\mathrm{MSE}_{s}}$ (5)
    is satisfied (with a regularization at the equator since both the meridional MSE transport and the meridional MSE gradient vanish). Here, the brackets denote vertical integration, the overline denotes a time and zonal average and the subscript “s” denotes surface quantities. The expression for the diffusivity coefficient comes from the hypothesis that the meridional energy flux is proportional to the gradient of the chosen generalized enthalpy at the surface (i.e. moist static energy in this case). In both cases, we use the same diffusivity for the control and superrotating runs. Taking into account, or not, the meridional structure in diffusivity does not affect the results.
@@
  • Moist model with varying relative humidity: Temperature changes satisfy $\delta R+\partial_{y}\left(D\partial_{y}\left(\delta\mathrm{MSE}/c_{p}\right)\right)=0$, where $\delta\mathrm{MSE}=c_{p}\delta T+L_{\rm v}\delta(\mathrm{RH}q^{\star}(T))$. The relative humidity profiles for the two states are shown in Figure 9 (panel b).
@@
We use the dedalus framework (dedalus) to solve for the equilibrium (i.e., steady state) temperature distribution as a boundary value problem, where we constrain the global average of temperature to be the same as in the GCM simulation and assume the meridional flux of energy vanishes at the poles. In the moist case, for the sake of simplicity, we linearize the temperature dependence of the saturation specific humidity, writing $q^{\star}(T)=\alpha T+\beta$.
@@ 6 Moisture transport and precipitation changes
It follows from the water mass budget that the $\mathrm{P-E}$ changes associated with a change in $F_{0}$ can be decomposed into a sum of different contributions, as done for example by seagerThermodynamicDynamicMechanisms2010; byrneResponsePrecipitationMinus2015 for anthropogenic climate change:
@@
Figure 10: Decomposition of $\mathrm{P-E}$ changes [see Eq. (6)] between the $F_{0}=0$ and $F_{0}=4$ m.s1-1.day1-1 experiments in the interactive- (panel a) and fixed-SST runs (panel b).
As shown in Figure 10, the different components of Eq. (6) behave differently when increasing the imposed torque. First of all, we note that both in the fixed- and interactive-SST experiments, two components of the decomposition are clearly more important than the others: the mean dynamic component and the eddy component. The two other components (mean thermodynamic and non-linear) are negligible. Note that this is qualitatively different from simulations of anthropogenic climate change where the thermodynamic effect tends to dominate, leading to an approximate Clausius-Clapeyron scaling for $\mathrm{P-E}$ (heldRobustResponsesHydrological2006), which holds relatively well over ocean regions (byrneResponsePrecipitationMinus2015).
We start by discussing the interpretation of the two dominant terms in the interactive-SST case (Fig. 10, a). The spatial structure of the total $\mathrm{P-E}$ change mostly matches the structure of the contribution due to MMC changes $\delta\mathrm{MDyn}$. These changes are directly related to the response of the MMC to the imposed tropical torque (Fig. 1) and are consistent with the meridional energy transport changes (Section 44.1): the collapse of the Hadley cells accounts for the negative sign of $\delta\mathrm{MDyn}$ within approximately $10^{\circ}$S-$10^{\circ}$N and its positive sign poleward of $12^{\circ}$ in each hemisphere. Similarly, the large shift of the Ferrell cells contributes to the poleward flank of the region of positive moisture convergence changes between $12^{\circ}$-$40^{\circ}$, and explains the negative changes around $50^{\circ}$ as well as the small positive contribution closer to the poles.
@@ 7 Conclusion
To unravel the mechanisms connecting the circulation changes to these temperature and precipitation changes, we have utilized a forcing/feedback analysis using radiative kernels along with decompositions of the energy and moisture transport changes into their various dynamical and thermodynamical components. These analyses reveal that different regions of the planet respond very differently to the imposed torque. First, the response in the tropics and subtropics is mostly governed by the moisture changes induced by the collapse of the MMC. This results in large relative humidity changes (decreasing in the tropics and increasing in the subtropics), which allow the atmosphere to compensate radiatively the changes in meridional energy transport related to circulation changes (larger energy convergence in the tropics, smaller in the subtropics). In other words, the radiative budget changes are dominated by the radiative forcing due to relative humidity, and feedbacks are small. Hence, in these regions surface temperature changes are also small. However, precipitation changes are large, and they could affect significantly the regional water cycle and therefore the biosphere. On the other hand, in the mid- to high-latitudes, the effects of moisture transport changes are too weak to compensate the energy transport changes, and the whole atmospheric column as well as the surface must warm to equilibrate the increased horizontal energy flux convergence. In other words, the temperature response is dominated by the Planck feedback. On a more technical note, this work demonstrates how insight can be gained into climate problems where the forcing is dynamic in origin, using various established conceptual tools (radiative forcing/feedback framework, fixed relative humidity ansatz, dry and moist EBMs, dynamic-thermodynamic decomposition of moisture transports).
Future work could investigate the robustness of the results presented here, for instance using different climate models and less idealized setups. In particular, the role of clouds or limited moisture availability at land surfaces are likely to influence the climate responses to forced circulation changes. In the case of clouds, previous results from Caballero2013 suggest that they could amplify the effects described here. It also remains to be better understood under which circumstances the climate of the Earth could spontaneously undergo such a transition, if at all, and how the two-way thermodynamic coupling between the processes driving the transition to superrotation and the impacts of superrotation described here would unfold. Finally, if the climate changes described here prove to be robust, a natural question would be whether they can be connected to paleo proxies for warm climates of the past.
@@ Acknowledgements.
Figure 11: TOA flux changes (panel a) across the transition to superrotation (from $F_{0}=0$ to $F_{0}=4$ m.s1-1.day1-1), and the decomposition of these changes into radiative forcings (panel b) and feedbacks (panel c) computed with radiative kernel. Panel a corresponds to the interactive-SST runs, panel b to the fixed-SST simulations and panel c is the difference between the two. Black curves represent the radiative budgets computed by the radiation scheme in the model, blue curves are the budgets reconstructed using the radiative kernels, which can be broken down into the Planck response (yellow curves), specific humidity response (red curves) and lapse-rate response (green curves). Numerical values given in the legend correspond to the global averages of the various components.
We start with the meridional structure of the radiative forcing (i.e., the TOA anomaly induced in the fixed-SST runs by prescribing $F_{0}=4$ m.s1-1.day1-1). We show in Figure 12 (panel b) and Figure 13 (panel b) the time- and zonally-averaged specific humidity and temperature changes, respectively. In the superrotating state, the atmosphere is drier than the control in the tropics, moister in the mid-latitudes, and slightly drier close to the poles. This explains qualitatively the meridional structure of the contribution of specific humidity changes to the radiative forcing (Fig. 11, panel b): drier air results in decreased optical depth, increased OLR, and negative radiative forcing, all else equal (conversely for moister air). Similarly, Figure 13 (panel b) shows that the lapse rate increases in the tropics, decreases in mid-latitudes, and increases again close to the poles, which matches the meridional structure of the lapse rate contribution to radiative forcing (Fig. 11, panel b).
Figure 12: Time- and zonally-averaged fractional specific humidity changes between the $F_{0}~=~4$ m.s1-1.day1-1 and $F_{0}~=~0$ m.s1-1.day1-1 experiments with interactive (panel a) and fixed SST (panel b). Here and for Figure 13, the dashed and solid black lines show the time- and zonally-averaged emission heights for $F_{0}=0$ and $F_{0}=4$ m.s1-1.day1-1, respectively.
Figure 13: Time- and zonally-averaged atmospheric temperature changes between the $F_{0}~=~4$ m.s1-1.day1-1 and $F_{0}~=~0$ m.s1-1.day1-1 experiments with interactive (panel a) and fixed SST (panel b).
However, the way the two effects combine into the net radiative forcing is more subtle (Fig. 11, panel b). To understand it, we consider jointly changes in the average emission level due to specific humidity changes and the spatial distribution of temperature changes (Figs. 12 and  13). Note that the emission level is defined as $p_{\rm em}(\varphi)=\left\{p(\varphi)\text{ such that }T(p,\varphi)=\left(\frac{\rm\overline{OLR}(\varphi)}{\sigma}\right)^{1/4}\right\}$, where $\sigma$ is the Stefan-Boltzmann constant. Close to the equator (below $12^{\circ}$ latitude), the drier atmosphere results in a substantially lower emission level, which is not compensated for by the lapse rate changes. Although the temperature in the superrotating state at the new emission level is cooler than it was at the same level in the control run, it is warmer than the temperature at the emission level in the control run, hence the OLR is larger: specific humidity effects dominate. Moving poleward, the competition between the two effects progressively changes: while the atmosphere is still drier in the superrotating state, it is less so and the emission level changes are smaller. At the same time, the lapse rate changes amplify. The combined change results in decreased OLR changes, until it vanishes around 12 latitude and even becomes negative when lapse rate effects dominate between $12^{\circ}$ and $20^{\circ}$. Between $20^{\circ}$ and $40^{\circ}$, the two effects reverse (Fig. 13): the lapse rate changes decrease until they become negative around $30^{\circ}$ latitude and poleward, and the atmosphere progressively becomes moister than the control run (Fig. 12), resulting in higher emission levels. As a result, the lapse rate contribution to the radiative forcing decreases with latitude in this region (Fig. 11, panel b) while the specific humidity contribution increases before levelling off. The fact that the OLR is smaller than in the control run in this region (the blue curve is positive) is therefore driven first by the specific humidity effects in the first half of the region, then by lapse rate effects in the second half. Going further poleward, the two effects essentially cancel out, resulting in a vanishing radiative forcing up to very high latitudes.
For the interactive-SST experiments, Figure 12 (panel a) shows that the specific humidity changes are smaller than in the fixed-SST case equatorward of $20^{\circ}$ (in this region, the emission level remains the same) — the superrotating state is actually moister except in the region where the OLR value is larger. Lapse rate changes are essentially smaller everywhere (Fig. 13, panel a). This explains the structure of the radiative feedback curves (Fig. 11, panel c), whose sign is opposite to the radiative forcing everywhere for the lapse rate component (green curve), and opposite in the tropics but the same sign in mid-latitudes for the specific humidity component (red curve; in other words specific humidity acts as a positive feedback everywhere). As a result, the specific humidity effect on changes in the net radiative budget vanishes in the tropics while the lapse rate component remains relatively small everywhere (Fig. 11, panel a). The total radiative budget changes are dominated by the Planck response in the deep tropics (with a minor contribution from the lapse rate), and become dominated by the imperfect cancellation between the Planck and specific humidity responses at higher latitudes.
Overall, the TOA radiative budget changes are consistent with changes in physical quantities in the atmosphere, such as specific humidity and temperature, but result from cancellations between the effects associated with these variables. As shown in section 44.2, these cancellations indicate that the effects of specific humidity and lapse rate on the radiative budget are essentially a consequence of the Clausius-Clapeyron relation. Hence, the radiative budget changes are better described by changes in relative humidity, which we use in the main text as a state variable instead of specific humidity.
[B]
[…]
Derivation of the relative humidity radiative response
We derive the expression for the relative humidity-based decomposition of TOA radiative budget changes (introduced in Section 44.2), starting from the specific humidity-based decomposition:
$\delta R$ $=\delta R(T_{\mathrm{s}},T,q)\approx\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}\delta T_{\mathrm{s}}+\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}\delta T+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}\delta q$
$\approx\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}\delta T_{\mathrm{s}}+\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}\delta T+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}\left(q^{\star}\delta\mathrm{RH}+\mathrm{RH}\frac{\partial q^{\star}}{\partial T}\delta T\right)$ (7)
$\approx\left(\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}\frac{\partial q}{\partial T}\right)\delta T+\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}\delta T_{\mathrm{s}}+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}q^{\star}\delta\mathrm{RH}.$
To split the atmospheric temperature change into a vertically-uniform component (the Planck response) and a vertically-resolved anomaly (the lapse rate response), we can write $\delta T=\delta T_{\mathrm{s}}+\delta(T-T_{\mathrm{s}})$ and we therefore have:
$\delta R\approx$ $\left(\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}\frac{\partial q}{\partial T}\right)\delta(T-T_{\mathrm{s}})$
$+\left(\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}\frac{\partial q}{\partial T}+\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}\right)\delta T_{\mathrm{s}}$ (8)
$+\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}q^{\star}\delta\mathrm{RH}.$
The radiative kernels are expressed as:
$\left\{\begin{array}[]{c}K^{T_{\mathrm{s}}}=\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}\\ K^{T}=\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}\\ K^{\mathrm{wv}}=\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\frac{\partial\log(q)}{\partial T}.\end{array}\right$. (9)
Since $\left.\frac{\partial R}{\partial q}\right|_{T_{\mathrm{s}},T}q^{\star}=\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\frac{1}{\mathrm{RH}}$, we can change the state variables and write:
$\begin{pmatrix}\delta R_{q}\\ \delta R_{T_{\mathrm{s}}}\\ \delta R_{\Gamma}\end{pmatrix}=\begin{pmatrix}K^{\mathrm{wv}}\frac{\delta\log(q)}{\partial\log(q)/\partial T}\\ \left(K^{T_{\mathrm{s}}}+K^{T}\right)\delta T_{\mathrm{s}}\\ K^{T}\left(\delta T-\delta T_{\mathrm{s}}\right)\end{pmatrix}\xrightarrow[\mathrm{{state\,variables}}]{\mathrm{Change\,of}}\begin{pmatrix}\widetilde{\delta R}_{\mathrm{RH}}\\ \widetilde{\delta R}_{T_{\mathrm{s}}}\\ \widetilde{\delta R}_{\Gamma}\end{pmatrix}=\begin{pmatrix}K^{\mathrm{wv}}\frac{\delta\log\left(\mathrm{RH}\right)}{\partial\log(q)/\partial T}\\ \left(K^{T_{\mathrm{s}}}+K^{T}+K^{\mathrm{wv}}\right)\delta T_{\mathrm{s}}\\ \left(K^{T}+K^{\mathrm{wv}}\right)\left(\delta T-\delta T_{\mathrm{s}}\right)\end{pmatrix}$, (10)
where the tildes indicate the different choice of independent variables (the tildes are not included in the main text).
We note that the relative-humidity component of $\widetilde{\delta R}$ can be written as:
$\widetilde{\delta R}_{\mathrm{RH}}=\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\delta\log(\mathrm{RH})$ $=\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\frac{1}{\mathrm{RH}}\frac{1}{q^{\star}}\left(\delta q-\mathrm{RH}\delta q^{\star}\right)$
$=K^{\mathrm{wv}}\frac{\delta\log(q)}{\partial\log(q)/\partial T}-\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\frac{1}{q^{\star}}\frac{\partial q^{\star}}{\partial T}\delta T$ (11)
$=K^{\mathrm{wv}}\frac{\delta\log(q)}{\partial\log(q)/\partial T}-K^{\mathrm{wv}}\delta T,$
which closes the description of the method used for the computations presented in Section 44.2.